Grade 12Trigonometry

if f(x) satisfies the relation 2f(x)+f(1-x)=x^2 for all real x, then find f(x).

Priya
8 Years agoGrade 12

1 Answer

suvidhi mehta
8 Years ago
2f(x)+f(1-x)=x^2 
this is a functional equality replace x by 1-x in whole equation
2f(1-x)+f(x)=(1-x)^2
now i think you can solve the above two equation to eliminate f(1-x) and get f(x)
i hope it helps you reach the answer